A random variable ℎ has a normal distribution with mean 7 and standard deviation
2.Calculate the probability that :
a)ℎ > 9
b)ℎ < 6
c)5 < ℎ < 8
1
Expert's answer
2021-10-08T04:05:53-0400

We are given,

"\mu=7"

"\sigma=2"

To find these probabilities, the values are standardized and the probabilities read from the standard normal tables as below.

a.

"p(h\gt9)=p(((h-\mu)/\sigma)\gt(9-\mu)/\sigma)=p(Z\gt(9-7)/2)=p(Z\gt1)"

This can also be written as,

"1-p(Z\lt1)=1-0.8413=0.1587"

Therefore "p(h\gt9)=0.1587"

b.

"p(h\lt6)=p(((h-\mu)/\sigma)\lt(6-\mu)/\sigma)=p(Z\lt(6-7)/2)=p(Z\lt-0.5)"

From the standard normal tables,

"p(Z\lt-0.5)=0.3085"

Hence, "p(h\lt6)=0.3085"

c.

"p(5\lt h\lt8)". On standardizing this we have,

"p((5-7)/2\lt Z\lt(8-7)/2)=p(-1\lt Z\lt 0.5)"

This probability can be written as,

"p(Z\lt0.5)-p(Z\lt-1)=0.6915-0.1587=0.5328"

Therefore, "p(5\lt h\lt 8)=0.5328" .